HIGH RESOLUTION CAVITY RING DOWN SPECTROSCOPY OF THE 3_{0}^{1} and 3_{0}^{1} 4_{0}^{1} BANDS OF THE Ã^{2}E′′ STATE OF NO_{3} RADICAL

\underline{MOURAD ROUDJANE}, TERRANCE J. CODD and TERRY A. MILLER, Laser Spectroscopy Facility, The Ohio State University, Columbus, Ohio 43210.

The NO_{3} radical is expected to exhibit a Jahn-Teller (JT) effect in its degenerate Ã^{2}E′′ electronic state.^{a} In the Ã state there are two JT active modes, ν_{3} and ν_{4} (e′ stretch and in-plane bend respectively). Theoretical work has predicted that the JT effect in the Ã state should be quite strong and approach the static case where the molecule is permanently distorted to a lower symmetry geometry. A more comprehensive understanding of its structure can be achieved using high resolution rotationally resolved absorption spectroscopy of its different vibronic bands. The high resolution absorption spectra of 3_{0}^{1} and 3_{0}^{1} 4_{0}^{1} vibronic bands of the Ã ^{2}E′′ excited state of NO_{3} have been successfully recorded for the first time using our jet cooled cavity ring down apparatus. These parallel bands are vibronically allowed transitions and shows the same contour. An oblate symmetric top model Hamiltonian including both centrifugal distortion and spin rotation terms is used to analyze the spectrum. The rotational analysis of this band, supported by combination differences, demonstrate the existence of doubled lines as were observed for 4_{0}^{n} (n=1,3) bands.^{b}

aE. Hirota, T. Ishiwata, K. Kawaguchi, M. Fujitake, N. Ohashi, and I. Tanaka, J. Chem. Phys., \underline{107}, 2829, 1997.

bRoudjane, M. et al. 67^{th} OSU International Symposium on Molecular Spectroscopy, The Ohio State University, Columbus, Ohio, 2011, talk TI-03.